### v1

# Can Big Black Holes Merge with the Smallest Black Holes?

Abstract
Gravitational-wave measurements of the binary black hole population provide insights into the evolution of merging binaries. We explore potential correlation between mass and mass ratio with phenomenological population models where the minimum mass of the smaller (secondary) black hole can change with the mass of the bigger (primary) black hole. We use binary black hole signals from the third Gravitational-Wave Transient Catalog with and without the relatively extreme mass-ratio GW190814. When excluding GW190814, models with a variable minimum mass are disfavoured compared to one with a constant minimum mass, with log Bayes factors of 2.49-2.49 to 0.98-0.98, indicating that the biggest black holes can merge with the smallest. When including GW190814, a parabola model that allows the minimum mass to decrease with increasing primary mass is favoured over a constant minimum-mass model with a log Bayes factor of $4.44$. When allowing the minimum mass to decrease, the overall population distributions remain similar whether or not GW190814 is included. This shows that with added model flexibility, we can reconcile potential outlier observations within our population. These investigations motivate further explorations of correlations between mass ratio and component masses in order to understand how evolutionary processes may leave an imprint on these distributions.
software: Bilby (Ashton et al., 2019), gwpopulation (Talbot et al., 2019), gwpopulation_pipe (Talbot, 2021), dynesty (Speagle, 2020), NumPy (Harris et al., 2020), matplotlib (Hunter, 2007), corner (Foreman-Mackey, 2016).
1 Introduction
[…]
GW observations of merging binary black holes from LIGO (Aasi et al., 2015), Virgo (Acernese et al., 2015), and KAGRA (Akutsu et al., 2021) allow us to probe the evolution of BBHs (Mapelli, 2020; Mandel & Farmer, 2022; Abbott et al., 2023a). The observed population likely contains contributions from multiple formation channels (Zevin et al., 2021; Cheng et al., 2023; Li et al., 2024; Colloms et al., 2025a), but the details of the astrophysical evolution of these systems are uncertain (Mandel & Broekgaarden, 2022; Belczynski et al., 2022; Mandel & Farmer, 2022). Measuring the underlying distribution of BBH masses (primary mass $m_{1}$ and secondary mass $m_{2}$), mass ratios ($q=m_{2}/m_{1}\leq 1$), spins, redshifts, and the correlations between these parameters from current gravitational-wave (GW) observations can inform us of formation mechanisms (Mapelli, 2021; Abbott et al., 2023a; Callister, 2024).
BBH formation processes are expected to leave signatures in the distribution of component masses and mass ratios of merging binaries. The mass ratios of binary stars during their evolution depends on the initial fragmentation, accretion during star formation and evolution, and supernovae mechanisms (Bate & Bonnell, 1997; Belczynski et al., 2012; Moe & Di Stefano, 2017; Neijssel et al., 2019; Zevin et al., 2020; Broekgaarden et al., 2022; Briel et al., 2022). Population-synthesis simulations indicate that field binaries that undergo a common-envelope stage can produce BBHs with mass-ratio distributions with support down to $q\sim 0.2$, though this depends on the treatment of the common-envelope process (Zevin et al., 2020; Olejak et al., 2021). Simulations of binaries that undergo stable mass transfer predict this channel cannot produce such unequal mass binaries, but can produce preferentially more massive black holes than common envelope (van Son et al., 2022). In contrast, BBHs that merge due to dynamical encounters may have more unequal masses and higher masses than isolated formation channels, especially if they are hierarchical mergers (Rodriguez et al., 2019; Sedda et al., 2023, 2020; Bruel et al., 2024). The von Zeipel–Lidov–Kozai effect for field triples can produce BBHs with mass ratios down to $0.3$ (Martinez et al., 2022). Similarly, active galactic nuclei can produce BBHs with more unequal masses than isolated binaries (Yang et al., 2019; Tagawa et al., 2020; Secunda et al., 2020). High-mass BBHs with unequal-masses could therefore be a signature of formation via dynamical encounters or more exotic processes.
Investigations of the mass distributions of black holes have found structure in the primary masses, secondary masses, and the mass ratios of the population. Fishbach et al. (2018) found that black holes do not randomly merge with each other regardless of mass; rather BBHs prefer equal masses. This is seen in the inferred mass-ratio distribution, which peaks at equal masses when considering a power-law model (Abbott et al., 2021, 2023a). Despite this preference for equal masses, secondary masses have been shown to follow a different distribution than primary masses, indicating that primary and secondary black holes undergo different astrophysical processes (Farah et al., 2024). Works considering more flexible models for the mass-ratio distribution find that the distribution peaks away from equal masses (Sadiq et al., 2024; Godfrey et al., 2023; Rinaldi et al., 2024; Callister & Farr, 2024a; Rinaldi et al., 2025). Furthermore, there is evidence that the mass-ratio distribution may change with primary mass (Sadiq et al., 2024) and that different features in the primary-mass distribution have different mass-ratio distributions (Li et al., 2022; Godfrey et al., 2023; Galaudage & Lamberts, 2025). These results indicate that there is information in the mass distributions that could be used to uncover the uncertain astrophysics of BBH formation (e.g., Colloms et al., 2025a). Exploring parametric model variations for BBH masses allows us to test assumptions about the relationship between component masses.
Callister (2024) highlights that the correlations between parameters, in addition to the parameters’ marginal distributions, can be informative. When investigating correlations, it is necessary to check whether results are accurate descriptions of the population, or whether they are driven by the enforced model constraints. Adamcewicz & Thrane (2022) investigated the anti-correlation between mass ratio and effective inspiral spin. By fixing the shape of the marginal distributions, they avoid the case where the correlation results from an improved fit of the marginal distributions. The same can be true in the other direction: inferences of marginal distributions could be driven by assumed correlations (or the lack thereof) imposed by the chosen model. Potentially, individual source constraints that are not well accounted for by model correlations could influence the measured marginal distributions, especially if a correlation must be distorted to fit them. Updating population models so that outliers are well accounted for and correlations are appropriately incorporated helps us understand the entire population of BBHs.
Modelling BBHs mass distributions requires a minimum black hole mass. This raises the question of the classification of merging objects with a mass that could correspond either to a black hole or a neutron star (Farah et al., 2022). It is unknown whether the maximum neutron star mass is the same as the maximum black hole mass. The maximum neutron star mass is uncertain, and depends on the neutron star equation of state (Rutherford et al., 2024; Koehn et al., 2025). Meanwhile, uncertainties in the minimum black hole mass come from the details of supernovae and core collapse (Fryer et al., 2012; Liu et al., 2021; Patton et al., 2022; Fryer et al., 2022). Following X-ray binary observations, it has been proposed that the maximum neutron star mass and the minimum black hole mass are not equal (Bailyn et al., 1998; Ozel et al., 2010; Farr et al., 2011; Siegel et al., 2023), and there is a mass gap with a dearth of black holes. Observations of objects with masses $\sim 2.5$$5~{M}_{\odot}$ indicate that at least some of the proposed gap is populated, though the rate of these objects is uncertain (Thompson et al., 2019; Abbott et al., 2020; Jayasinghe et al., 2021; Abbott et al., 2023b; Abac et al., 2024; Fishbach et al., 2025). Including low-mass black holes from GW observations can allow us to measure the minimum black hole mass (Ray et al., 2025).
Observations of (potential) low-$m_{2}$ BBHs include GW190814, which has $q=0.112^{+0.008}_{-0.009}$ and $m_{2}=2.6^{+0.1}_{-0.1}{M}_{\odot}$ (Abbott et al., 2020, 2024). Understanding the nature of GW190814’s source as a BBH or a neutron star–black hole requires understanding if it is an outlier to the BBH population, and whether its outlier status is because of population misspecification. While Abbott et al. (2021) and Essick et al. (2022) find GW190814 to be an outlier, Farah et al. (2022) consider models where it is not by including the full population of neutron stars and black holes, and considering a partially full mass gap. The outlier status of observations thus depends on the choice of population model. We seek to understand how to improve our models and incorporate observations that are currently challenging to fit.
Current models for the BBHs mass distribution assume that the minimum black hole mass is constant with the primary black hole mass. Given the preference for equal mass mergers and the structure in the distribution of BBH masses, this may not be the case. We investigate if there is evidence for an evolving minimum secondary mass with primary mass: can black holes with higher masses merge with the lowest-mass black holes? We introduce models for the mass-ratio distribution of BBHs specifying a minimum secondary black hole mass that evolves with the primary mass (Section 2). Using these models on observations from the third Gravitational-Wave Transient Catalog (GWTC-3.0; Abbott et al., 2023b) we infer the primary-mass and mass-ratio distributions (Section 3). Our conclusions for the inferred evolution of minimum secondary mass, and measured primary-mass and mass-ratio distributions are discussed in Section 4.
2 Methods
We estimate a set of hyperparameters $\Lambda$ that govern the shape of the mass, spin and redshift distributions of GW observations from third Gravitational-Wave Transient Catalog (GWTC-3.0) (Abbott et al., 2023b) using a hierarchical Bayesian inference (Mandel et al., 2019; Thrane & Talbot, 2019; Vitale et al., 2020). For each population distribution we choose a parametric model. We assume the distribution of the primary mass, spin and redshift distributions to be the default parametric models used in Abbott et al. (2023a), and introduce new models for the mass-ratio distribution as described in Section 2.1. Our inference framework is described in more detail in section Section 2.2 and the observations we use are described in Section 2.3.
2.1 Variable minimum-mass models
To investigate the correlation between the minimum mass and the primary mass, we model the mass-ratio distribution of BBHs with a minimum secondary mass that varies with the primary mass. We include this as a modification to a power-law model of mass ratio (Fishbach & Holz, 2017; Kovetz et al., 2017), such that the mass-ratio distribution ${\mathrm{p}}(q\mid m_{1})$ with a power-law index $\beta$ and a minimum secondary mass $m_{2}^{\mathrm{min}}(m_{1})$ follows
$\begin{split}{\mathrm{p}}(q\mid m_{1})&=\left(1+\beta\right)\left[1-\frac{m_{2}^{\mathrm{min}}(m_{1})}{m_{1}}\right]^{-(1+\beta)}q^{\beta}\\ &\propto q^{\beta}\qquad\qquad\left\{\frac{m_{2}^{\mathrm{min}}(m_{1})}{m_{1}}\leq q\leq 1\right\}.\end{split}$ (1)
This allows a varying lower bound on the secondary mass $m_{2}$ such that $m_{2}\geq m_{2}^{\mathrm{min}}(m_{1})$: the smallest possible secondary mass, and therefore the minimum mass ratio varies with the primary mass.
We consider two forms for how the minimum secondary mass varies with primary mass. Neither has significant astrophysical motivation, but both are simple progressions from a fixed minimum mass that enable us to explore whether the data support more complicated models. The two models complement each other in their analytic forms, and minimise additional model hyperparameters.
First, we consider a Power Law form that adds one additional hyperparameter:
$m_{2}^{\mathrm{min}}=m_{\mathrm{min}}+(m_{\mathrm{max}}-m_{\mathrm{min}})\left(\frac{m_{1}-m_{\mathrm{min}}}{m_{\mathrm{max}}-m_{\mathrm{min}}}\right)^{\gamma},$ (2)
where $\gamma\geq 1$, $m_{\mathrm{min}}$ is the smallest primary black hole mass, and $m_{\mathrm{max}}$ is the maximum (stellar-mass) black hole mass. The Power Law model is a strictly increasing function that sets $m_{\mathrm{min}}$ as the minimum black hole mass for both the primary and secondary objects, and $m_{2}^{\mathrm{min}}(m_{\mathrm{max}})=m_{\mathrm{max}}$. As $\gamma$ increases, lower $m_{2}^{\mathrm{min}}$ is allowed for each $m_{1}$. Inspired by observations of binaries preferring more equal-mass components, this model prevents high-mass primaries from merging with low-mass secondaries.
Second, we consider a Parabola model for $m_{2}^{\mathrm{min}}$ that adds two additional hyperparameters:
$m_{2}^{\mathrm{min}}(m_{1})=m_{\mathrm{min}}+\xi(m_{1}-m_{\mathrm{min}})+\zeta(m_{1}-m_{\mathrm{min}})^{2}.$ (3)
This functional form automatically enforces that $m_{2}^{\mathrm{min}}(m_{\mathrm{min}})=m_{\mathrm{min}}$. We enforce additional constraints to ensure that the model remains physical. We require that $m_{2}^{\mathrm{min}}(m_{1})<m_{1}$ for all $m_{1}<m_{\mathrm{max}}$. Equivalently,
$m_{\mathrm{max}}>m_{\mathrm{min}}+\xi(m_{\mathrm{max}}-m_{\mathrm{min}})$
$+\left.\zeta(m_{\mathrm{max}}-m_{\mathrm{min}})^{2}.\right$. (4)
This yields the bound
$\zeta<\frac{1-\xi}{m_{\mathrm{max}}-m_{\mathrm{min}}}.$ (5)
We consider two variations of this model.
The Increasing Parabola model has $\xi>0$ and $\zeta>0$ such that $m_{2}^{\mathrm{min}}$ is always increasing with $m_{1}$.
The Relaxed Parabola model allows for a decreasing $m_{2}^{\mathrm{min}}$ with $m_{1}$ by allowing for negative $\xi$. This means that the minimum secondary mass can be less than $m_{\mathrm{min}}$. For this case, we require explicitly enforcing that $m_{2}^{\mathrm{min}}(m_{1})>0$ for all $m_{1}$. This necessitates that Eq. (3) has no real roots, and
$\zeta>\frac{\xi^{2}}{4m_{\mathrm{min}}}.$ (6)
Combining Eq. (5) and Eq. (6) sets a limit on $\xi$ such that
$\frac{1-\xi}{m_{\mathrm{max}}-m_{\mathrm{min}}}>\frac{\xi^{2}}{4m_{\mathrm{min}}},$ (7)
and hence,
$-2\frac{\sqrt{m_{\mathrm{min}}}}{\sqrt{m_{\mathrm{max}}}-\sqrt{m_{\mathrm{min}}}}<\xi<2\frac{\sqrt{m_{\mathrm{min}}}}{\sqrt{m_{\mathrm{max}}}+\sqrt{m_{\mathrm{min}}}}.$ (8)
We consider these constraints in our priors on $\xi$ and $\zeta$, as detailed in Table 1.
Model Hyperparameter priors Minimum secondary mass
Power Law $\gamma\sim\mathcal{U}(1,35)$ Eq. (2)
Increasing Parabola $\xi\sim\mathcal{U}(0,1)$ Eq. (3)
$\zeta\sim\mathcal{U}\left(0,\frac{1-\xi}{m_{\mathrm{max}}-m_{\mathrm{min}}}\right)$
Relaxed Parabola $\xi\sim\mathcal{U}\left(-2\frac{\sqrt{m_{\mathrm{min}}}}{\sqrt{m_{\mathrm{max}}}-\sqrt{m_{\mathrm{min}}}},2\frac{\sqrt{m_{\mathrm{min}}}}{\sqrt{m_{\mathrm{max}}}+\sqrt{m_{\mathrm{min}}}}\right)$ Eq. (3)
$\zeta\sim\mathcal{U}\left(\frac{\xi^{2}}{4m_{\mathrm{min}}},\frac{1-\xi}{m_{\mathrm{max}}-m_{\mathrm{min}}}\right)$
Table 1: Prior ranges on $\gamma$, $\xi$ and $\zeta$ hyperparameters for the three minimum secondary-mass models we consider, Power Law, Increasing Parabola, and Relaxed Parabola, and their corresponding forms for the minimum secondary mass.
2.2 Statistical framework
To investigate the overall distribution of the BBH population, we jointly infer the distribution of BBH primary masses, mass ratios, spin magnitudes and tilts, and redshifts. We model the distribution of BBH primary masses, mass ratios, spin magnitudes $\chi_{1,2}$ and tilts $\theta_{1,2}$, and redshifts $z$ as a separable probability distribution conditional on population hyperparameters $\Lambda$,
$\begin{gathered}p(m_{1},q,\chi_{1,2},\cos\theta_{1,2},z\mid\Lambda)=\left.p(m_{1}\mid\Lambda)\times p(q|m_{1},\Lambda)\right.\times\\ p(\chi_{1,2}\mid\Lambda)\times p(\cos\theta_{1,2}\mid\Lambda)\times p(z\mid\Lambda),\end{gathered}$ (9)
where the mass-ratio distribution includes the Power Law, Increasing Parabola or Relaxed Parabola minimum secondary-mass model described above, and the other observables have distributions that follow the default parametric models used in Abbott et al. (2023a). We use the Power Law+Peak model for $p(m_{1})$ (Talbot & Thrane, 2018). This models the $m_{1}$ distribution as a two component mixture of a power-law of slope $\alpha$ and a Gaussian component. We choose the similar priors on the default model parameters as Abbott et al. (2023a), with some refinements. We use primary-mass power-law index $\alpha\sim\mathcal{U}(-4,12)$, mass-ratio power-law index $\beta\sim\mathcal{U}(-2,7)$, primary minimum mass $m_{\mathrm{min}}\sim\mathcal{U}(2,6)\,M_{\odot}$, and maximum mass $m_{\mathrm{max}}\sim\mathcal{U}(62,100)\,M_{\odot}$. Both spin magnitudes are modelled with an identical beta distribution, and the spin tilts a mixture of isotropic and normally distributed spin tilts. The redshift distribution $p(z)$ is modelled with a power law (Fishbach et al., 2018) of index $\kappa\sim\mathcal{U}(-2,8)$ .
To infer hyperposteriors on $\Lambda$ we use the population-inference package gwpopulation_pipe (Talbot, 2021) with the nested sampler dynesty (Speagle, 2020). The inference imposes a cut on the likelihood variance of 11 during sampling (Talbot & Golomb, 2023).
2.3 Gravitational-wave observations
We use the observations from third Gravitational-Wave Transient Catalog (GWTC-3.0; Abbott et al., 2023b) selected with a false alarm rate threshold of less than $1~\mathrm{yr}^{-1}$. This is identical to the set of observations used in Abbott et al. (2023a). We consider only signals with BBH sources and the ambiguous GW190814 (Abbott et al., 2020, 2024). As GW190814’s secondary object may be above the expected maximum mass for a neutron star, we compare the results with and without GW190814, and investigate its potential as an outlier to the BBH population. GW200210_092254 (Abbott et al., 2023b), which has a similar secondary mass and mass ratio to GW190814, is not included in our analysis as it does not satisfy the false alarm rate threshold.
We use a combined injection set of semi-analytic injections at the sensitivity of the first and second observing runs, and real injections at the third observing run sensitivity to account for selection effects (Abbott et al., 2023a). We consider the real injections with a false alarm rate of less than $1~\mathrm{yr}^{-1}$ detectable, and the semi-analytic injections we cut in network signal-to-noise ratio of $10$ (Abbott et al., 2018).
3 Results
3.1 Evolution of minimum secondary mass
Figure 1: Inferred evolution of the minimum secondary mass with primary mass for the Power Law (left), Increasing Parabola (middle), and Relaxed Parabola (right) models. Each panel shows the resulting median and $90\%$ credible interval on the minimum secondary mass, including and excluding GW190814. The grey shows the $99\%$ credible interval for the prior for each model, with the lower edge in black. The interval is evaluated over $m_{1}\in[2,100]~M_{\odot}$, where the plotted range extends to $m_{1}>m_{\mathrm{max}}$ for some individual hyperposterior samples. With GW190814, the entire distribution of $m_{2}^{\mathrm{min}}$ is lower, apart from the Relaxed Parabola model, which decreases to allow for the lower secondary mass at $m_{1}\sim 20\,M_{\odot}$. Due to the driving of the prior, $m_{2}^{\mathrm{min}}$ then increases at higher $m_{1}$ for this model. Otherwise, there is no strong evidence for an increasing $m_{2}^{\mathrm{min}}$ with $m_{1}$ when accounting for model limitations.
We show the inferred evolution of $m_{2}^{\mathrm{min}}$ with $m_{1}$ in Figure 1, with and without GW190814, for our three model variations. Uncertainties grow at larger $m_{1}$ where we have fewer observations.
Under the Power Law minimum secondary-mass model, the evolution of $m_{2}^{\mathrm{min}}$ is preferentially flat until high primary masses, where the uncertainty on $m_{2}^{\mathrm{min}}$ grows. At high $m_{1}$, $m_{2}^{\mathrm{min}}$ is forced to increase to $m_{\rm{max}}$ because of the functional form of the model. The hyperposteriors in Figure 2 show that high $\gamma$ is preferred, meaning that the evolution of $m_{2}^{\mathrm{min}}$ is flatter with $m_{1}$, before sharply increasing as $m_{1}$ approaches $m_{\mathrm{max}}$. The results are restricted by the prior on $\gamma$, as we allow a maximum $\gamma$ of $35$, while a flat evolution in $m_{2}^{\mathrm{min}}$ would require $\gamma\rightarrow\infty$. With our prior choice, we find that $\gamma>4.4$ with GW190814 and $\gamma>2.8$ without GW190814 at $90\%$ credibility. The preference for a flatter evolution of $m_{2}^{\mathrm{min}}$ means that there is little support for a higher minimum secondary mass at higher primary masses. While this behaviour is consistent whether or not GW190814 is included, the overall minimum mass $m_{\mathrm{min}}$ and the primary mass power-law index $\alpha$ are different when including GW190814. Without GW190814, the minimum mass is higher, with $m_{\mathrm{min}}=5.8^{+0.2}_{-0.6}$, while with GW190814, $m_{\mathrm{min}}=2.2^{+0.4}_{-0.2}$ to accommodate GW190814’s secondary mass.
Figure 2: Hyperposteriors on the key mass and mass-ratio hyperparameters with Power Law model, with and without GW190814. The $m_{2}^{\mathrm{min}}$ hyperparameter, $\gamma$, prefers high values, supporting a non-increasing or shallowly increasing $m_{2}^{\mathrm{min}}$ with $m_{1}$. The other hyperparameters show disagreement between the choice to include or exclude GW190814.
For the Increasing Parabola model, the inferred $m_{2}^{\mathrm{min}}$ behaviour is similar to the Power Law result, where $m_{2}^{\mathrm{min}}$ is consistent with being flat with $m_{1}$ for both sets of results. Hence, there is no evidence that the minimum secondary mass changes with primary mass when considering this model. This can be seen in the hyperposteriors on $\xi$ and $\zeta$, as plotted in Figure 3, with $\xi$ and $\zeta$ both peaking at 0. Similarly to the Power Law model, when including GW190814, $m_{\mathrm{min}}$ is forced to be lower to include its secondary mass, and $\alpha$ is smaller. At large primary masses, the uncertainty on the minimum secondary mass is high, though this is reduced when including GW190814. This may result from the restriction that $m_{2}^{\mathrm{min}}$ cannot increase above $\sim 2.6~M_{\odot}$ at $m_{1}\sim 20~M_{\odot}$, as can be seen by the more tightly constrained $\xi$ and $\zeta$ when including GW190814.
Figure 3: Hyperposteriors on the key mass and mass-ratio hyperparameters with the Increasing Parabola model, with and without GW190814. The $m_{2}^{\mathrm{min}}$ hyperparameters, $\xi$ and $\zeta$ are railing at 0, showing a preference for a non-increasing $m_{2}^{\mathrm{min}}$ with $m_{1}$. The other hyperparameters show disagreement between the choice to include or exclude GW190814.
In the Relaxed Parabola model, $m_{2}^{\mathrm{min}}$ remains largely unconstrained at high $m_{1}$, while $m_{2}^{\mathrm{min}}(m_{\mathrm{max}})=m_{\mathrm{min}}$ is not ruled out when not including GW190814. With GW190814, $m_{2}^{\mathrm{min}}$ decreases below $m_{\mathrm{min}}$ until $m_{1}\sim 20~M_{\odot}$, then increases at high $m_{1}$, preferring $m_{2}^{\mathrm{min}}>m_{\mathrm{min}}$ at higher primary masses. The dip in $m_{2}^{\mathrm{min}}$ at $m_{1}\sim 20~M_{\odot}$ accommodates GW190814’s secondary mass, while the increase of $m_{2}^{\mathrm{min}}$ at high $m_{1}$ is driven by the prior constraints when including negative $\xi$. For both sets of observations, $m_{\mathrm{min}}$ does not significantly change as $m_{2}^{\mathrm{min}}$ is instead allowed to decrease to allow for the secondary mass of GW190814. This is because $m_{\mathrm{min}}$ is no longer describing the lowest secondary mass of the observations, but the lowest primary mass. As shown in Figure 4, while the hyperposteriors on $\xi$ and $\zeta$ are different with the different sets of observations, the other hyperposteriors show less variation when including GW190814 than results with the strictly increasing $m_{2}^{\mathrm{min}}$ models. This hints that GW190814 is no longer an outlier under the model assumptions that allow $m_{2}^{\mathrm{min}}$ to decrease with $m_{1}$.
Figure 4: Hyperposteriors on the key mass and mass ratio hyperparameters with Relaxed Parabola model, with and without GW190814. The $m_{2}^{\mathrm{min}}$ hyperparameters, $\xi$ and $\zeta$ are consistent with zero without GW190814, but negative $\xi$ is preferred when including GW190814. The other hyperparameters are in agreement between the two choices of observations.
We compare each of our analyses to the results with the default Power Law + Peak model from Abbott et al. (2023a), with $m_{2}^{\mathrm{min}}$ fixed to be equal to $m_{\mathrm{min}}$. We find that our models are generally not preferred to the default model without a variable $m_{2}^{\mathrm{min}}$, as shown in Table 2. This reinforces the result that there is no evidence that $m_{2}^{\mathrm{min}}$ (strictly) increases with $m_{1}$. The exception is the Relaxed Parabola model, which is preferred with a log Bayes factor of $\log_{10}\mathcal{B}=4.44$ when including GW190814. Therefore, if GW190814’s source is a BBH, this indicates that there is structure in the mass-ratio and component-mass distributions that should be accounted for by population models.
$\log_{10}\mathcal{B}$
With Without
Model GW190814 GW190814
Power Law 0.54-0.54 0.98-0.98
Increasing Parabola 1.86-1.86 2.49-2.49
Relaxed Parabola $\hphantom{-}4.44$ 1.37-1.37
Table 2: Log Bayes factors between the Power Law, Increasing Parabola, and Relaxed Parabola minimum secondary-mass models and the default model without evolution of the minimum secondary mass from Abbott et al. (2023a). All models allowing for an evolving minimum secondary mass with primary mass are disfavoured compared to the default model, with the exception of the Relaxed Parabola model when including GW190814.
3.2 Mass-ratio and primary-mass distributions
Figure 5: The inferred primary-mass (top) and mass-ratio (bottom) distributions for the Power Law (left), Increasing Parabola (middle), and Relaxed Parabola (right) models, each with and without GW190814. The solid line shows the posterior population distribution and the shaded region is the $90\%$ credible interval for each distribution. For the Increasing Parabola and Power Law models, the distributions are different when including GW190814, while for the Relaxed Parabola model the intervals are consistent.
To understand the impact of a variable $m_{2}^{\mathrm{min}}$ on the mass-ratio distribution, we show $p(m_{1})$ and $p(q)$ for our three model variations in Figure 5. For the Power Law and Increasing Parabola models, neither $p(m_{1})$ nor $p(q)$ agree between the results with and without GW190814. Both of these models find a different $m_{\mathrm{min}}$ for the different sets of observations, resulting in a different $p(m_{1})$ distribution at lower masses. Without GW190814, these models also are not required to go to as low mass ratios. With these strictly increasing $m_{2}^{\mathrm{min}}$ models, inclusion of GW190814 changes the overall primary-mass distribution despite its $m_{1}$ being unexceptional. For the Relaxed Parabola model, there is more agreement between both primary-mass and mass-ratio distributions. While $p(q)$ does extend to lower values with GW190814, this is expected due to its low mass ratio. The added flexibility in the mass-ratio model thus allows for GW190814-like systems to fit with the bulk BBHs population, and they do not need to be treated as population outliers.
Figure 6 shows $p(q)$ at slices of $m_{1}$ for the three model variations. This shows the strictly increasing $m_{2}^{\mathrm{min}}$ models’ inconsistency between the mass-ratio distributions for the results with and without GW190814, especially for lower primary masses. At $m_{1}=10~M_{\odot}$ for the Power Law and Increasing Parabola models, the mass-ratio distribution supports much lower mass ratios when including GW190814 because of the overall lower $m_{\mathrm{min}}$. As $m_{\mathrm{min}}\sim 5~M_{\odot}$ without GW190814, the lowest possible mass ratio is $q\sim 0.5$ at $m_{1}=10~M_{\odot}$, while when including GW190814, $m_{\mathrm{min}}\sim 2~M_{\odot}$, making the lowest possible mass ratio $q\sim 0.2$. Extreme mass ratios at $m_{1}=10~M_{\odot}$ are not well supported by the prior, and GW190814 is highly informative for the inference. At all mass slices, $p(q)$ is flatter for inference with GW190814, driven by the available model space. The difference at $m_{1}=10~M_{\odot}$ is reconciled for the Relaxed Parabola model, as $m_{\mathrm{min}}$ is in agreement whether or not GW190814 is included.
Figure 6: Slices of the median and $90\%$ interval for the mass-ratio distribution at primary masses of $10~M_{\odot}$ (left), $20~M_{\odot}$ (middle), and $40~M_{\odot}$ (right) for the Power Law (top), Increasing Parabola (middle), and Relaxed Parabola (bottom) models, each with and without GW190814. For the Increasing Parabola and Power Law models, the distributions are different when including GW190814, while for the Relaxed Parabola model the intervals are more consistent.
Compared to our analyses, Callister & Farr (2024b) and Sadiq et al. (2024) find that $p(q)$ has more support for unequal-mass mergers when considering data-driven models. These models do not consider a parametrised $m_{\mathrm{min}}$ and therefore have more flexibility to extend support to lower $q$ even at low $m_{1}$. Sadiq et al. (2024), when excluding GW190814, find that $p(q\mid m_{1}=10~M_{\odot})$ has support down to $q\sim 0.3$, while the distribution peaks away from equal masses at higher $m_{1}$.
4 Conclusions
We investigated the correlation between mass ratio and primary mass of merging BBHs to understand whether the biggest stellar-mass black holes can merge with the smallest. We introduced population models that included a minimum secondary mass $m_{2}^{\mathrm{min}}(m_{1})$ that varies with primary mass. We considered three model variations for $m_{2}^{\mathrm{min}}$, which we used with a power-law model for the mass-ratio distribution and a Power Law + Peak distribution for primary mass (Talbot & Thrane, 2018; Abbott et al., 2023a). Using these models we inferred how the minimum secondary mass evolves for GWTC-3.0 observations (Abbott et al., 2023b).
Our results show that there is currently no evidence that $m_{2}^{\mathrm{min}}$ increases with $m_{1}$ when excluding GW190814. We find log Bayes factors of $\log_{10}\mathcal{B}=-2.49$ to 0.98-0.98 for our models compared to a constant minimum-mass model. We repeated the same analysis including GW190814. If we enforce that $m_{2}^{\mathrm{min}}$ is strictly increasing with $m_{1}$, we find that the overall minimum black hole mass has to be lower to account for GW190814’s secondary mass, while the primary-mass and mass-ratio distributions change with inclusion of GW190814. This suggests that GW190814 should be considered as an outlier for these population models. With a model allowing $m_{2}^{\mathrm{min}}$ to decrease with $m_{1}$, $m_{2}^{\mathrm{min}}$ dips to allow GW190814’s secondary mass, and the minimum primary mass is in agreement with the inference excluding GW190814. The added flexibility in $m_{2}^{\mathrm{min}}$ means that under this model the differences in $p(m_{1})$ and $p(q)$ are reconciled when doing a leave-one-out analysis, no longer causing GW190814 to be an outlier. This model is favoured when including GW190814 in the population compared to a constant minimum-mass model, with $\log_{10}\mathcal{B}=4.44$. Hence, there is support for an evolving $m_{2}^{\mathrm{min}}$ when including GW190814’s source in the BBH population.
The preference for a near-constant $m_{2}^{\mathrm{min}}$ with $m_{1}$ means that we do not rule out the possibility of high-primary mass, relatively extreme mass-ratio BBHs. However, our model uncertainties are greatest at large primary masses due to a combination of a lack of observations and model limitations.
While our models provide one way to understand correlations between primary mass and mass ratio, improvements to our model could be made to ensure robustness and avoid model misspecification. Our model that allows for decreasing $m_{2}^{\mathrm{min}}$ enforces that the minimum secondary mass is always positive, but currently does not place other bounds on $m_{2}^{\mathrm{min}}$, e.g., it could go to subsolar masses. Future work could update this assumption by imposing a non-zero lower bound on $m_{2}^{\mathrm{min}}$, representing a minimum black hole mass, which may help us understand the classification of GW190814. Additional model flexibility also could be added, for example by relaxing the assumption that $m_{2}^{\mathrm{min}}(m_{\mathrm{min}})=m_{\mathrm{min}}$; Farah et al. (2024) find that the inferred minimum primary and minimum secondary masses of the BBH population are in agreement within current uncertainties, but did not include GW190814 in their analysis. Our parametric forms for the minimum secondary mass are also not motivated by particular astrophysics or trend in the data. A data-driven description of the primary-mass and mass-ratio distributions may give more insight into the underlying correlations, unrestricted by the choice of parametric form. Furthermore, we specify a power-law distribution in mass ratio, restricting the mass-ratio distribution to be strictly monotonic. However, models with more relaxed assumptions about the shape of the distribution show support for the mass-ratio distribution to peak away from $1$ (Sadiq et al., 2024; Godfrey et al., 2023; Rinaldi et al., 2024; Callister & Farr, 2024a; Rinaldi et al., 2025). Considering a mass-ratio distribution that more accurately reflects the data would help avoid biases driven by an inaccurate models.
Other observable parameters such as the spin or eccentricity could provide more information about the astrophysical nature of such BBHs. Works looking at the correlations between mass and spin may identify subpopulations of sources from similar evolutionary pathways (e.g., Safarzadeh et al., 2020; Heinzel et al., 2024; Ray et al., 2024; Antonini et al., 2025). If the mass-ratio distribution is jointly correlated with component mass and spin distributions, neglecting to model these correlations could lead to biased inference (Alvarez-Lopez et al., 2025). Jointly investigating mass, mass-ratio, and spin distributions could help us understand how these parameters are correlated, and therefore give more insight into the astrophysical evolution of the population. However, adding additional parameters to our models and investigating the joint correlations between mass, mass ratio, and spin, requires having sufficient data to constrain the model parameters. The increased number of BBH observations coming with the forth LIGO–Virgo–KAGRA observing run will allow us to better measure any correlations between the BBH observables, and explore the astrophysics encoded in these correlations.
acknowledgments
We thank the groups at Northwestern University for hospitality while this work was developed, in particular to Sylvia Biscoveanu, Monica Gallegos-Garcia, and Miguel Martinez for insightful discussions. We thank Alexandra Guerrero for comprehensive review, Daniel Holz for valuable insights, and John Veitch for useful technical discussions. SC is supported by Science and Technology Facilities Council (STFC) studentship 2748218. CPLB is supported by STFC grant ST/V005634/1. ZD acknowledges support from the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA) Board of Visitors Research Professorship. SC thanks the CIERA Board of Visitors and the University of Glasgow MacRobertson Scholarship for support to visit CIERA. This research has made use of data from the Gravitational Wave Open Science Center, a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by the National Science Foundation (NSF) LIGO Laboratory which is a major facility fully funded by the NSF, as well as STFC of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO 600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by NSF Grants PHY-0757058 and PHY-0823459. This document has been assigned LIGO document number LIGO-P2500498.
These results were produced using models found in this code release. Results and data products used in this work are available on Zenodo: https://doi.org/10.5281/zenodo.16895784 (catalog doi:10.5281/zenodo.16895784) (Colloms et al., 2025b).